OpenStudy (anonymous):

Can someone explain how I would go about in completing this problem? If a segment is divided into two parts having lengths in the ratio of 5:3. If the difference between the length of the parts is 6", find the length of the longest part. A) 3 B) 9 C) 15 D) 24

OpenStudy (neer2890):

Let the lengths are 5x and 3x respectively. so According to Question: 5x-3x=6 2x=6 x=3 so, length of longer part=5x=5*3=15"

OpenStudy (anonymous):

Oh...I see now! Thanks man!

OpenStudy (neer2890):

you're welcome..:)